OpenStudy (anonymous):

What is the general term equation, an, for the arithmetic sequence 9, 5, 1, −3, . . . , and what is the 21st term of this sequence?

OpenStudy (bobo-i-bo):

Can you see what the pattern is?

OpenStudy (bobo-i-bo):

As in, firstly, can you describe the pattern which the arithmetic sequence follows?

OpenStudy (anonymous):

Hint: an = a1 + d(n − 1), where a1 is the first term and d is the common difference.

jhonyy9 (jhonyy9):

@chris5122 than you see it 5 is smaler than 9 by 4 - yes ? and so the other 1 is smaler than 6 by 4 - yes ? so how you think it ?

jhonyy9 (jhonyy9):

sorry for my mistake so 1 is smaler than 5 by 4

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

so than what will be the 21th term ?

OpenStudy (anonymous):

Thats what im asking help with..

OpenStudy (mathmale):

You're looking for the "common difference." From one term to the next, you see a decrease. How much is this decrease? Call it " d ." This is the "d" to which Chris is referring in : an = a1 + d(n − 1)

OpenStudy (anonymous):

Im really just looking for the 21st term

OpenStudy (mathmale):

Excuse me for saying so, Chris, but you're completely missing the point. To use the formula an = a1 + d(n − 1), you MUST know what "common difference" means and be able to find it. Were I to give you the 21st term, you'd learn next to nothing .

OpenStudy (mathmale):

What is the common difference? What is the jump / leap from 9 to 5, from 5 to 1, etc.?

OpenStudy (anonymous):

ohhhhh

OpenStudy (anonymous):

the pattern

OpenStudy (mathmale):

Yes, the pattern. ??

OpenStudy (anonymous):

so -4

OpenStudy (mathmale):

Yes, the "common difference" is d = -4. So much better. Now, write out that formula, replacing "d" with -4. an = a1 + d(n − 1)

OpenStudy (anonymous):

I think i may know it

OpenStudy (anonymous):

would it be an=9-4n, a21=-75

OpenStudy (mathmale):

the first term is 9. Call this "a1" ("a sub 1"). d = -4. To find the 21st term, subst. 21 for n in the equation.

OpenStudy (mathmale):

Please show how you got -75. Note that a1=9, d=-4, and n=21.

OpenStudy (anonymous):

one second im doing the work

OpenStudy (mathmale):

Excuse me, but where are you?

OpenStudy (anonymous):

Here

OpenStudy (anonymous):

Im sorry

OpenStudy (mathmale):

Your work, please?

OpenStudy (anonymous):

well since 9 was the first term i could figure out by process of elimination what the answer was. There is only one answer where the first term is 9.

OpenStudy (anonymous):

Its a multiple choice question so since a1=9 that gave me the answer.

OpenStudy (mathmale):

why not learn how to do this problem properly? Please use the equation\[l=a + (n-1)d\] where a= first term, d=common difference, n=the index of the term wanted, and l=last term. You 'll need to calculate the last (the 21st) term, d.

OpenStudy (mathmale):

a=? d=? n=? l=?

OpenStudy (anonymous):

Well better off doing it the right way..

OpenStudy (anonymous):

one second ill answer it

OpenStudy (anonymous):

I get confused when you say index of the term wanted?

OpenStudy (anonymous):

what do you mean by that?

OpenStudy (mathmale):

n is a counter, like your fingers are counters when you count 1, 2, 3, ... First term could be 

OpenStudy (mathmale):

2nd term could be